Brianna had $800 to invest last year She invested some in a garanteed 8% intrest account she put the rest in a riskier account and lost 4% on the investment if by the end of the year she had made $28 as a result of both investments how much did she invest in each account
Answer by rfer(16322)

x=guarnteed accout
800-x=riskier account
.08x-.04(800-x)=28
.08x-32+.04x=28
.12x=60
x=500
800-x=300
500*.08=40
300*.04=12
40-12=28